The average return of Intermediate-term government bonds is 5.4% per year and the standard deviation is 3.2%. What range of returns would you expect to see 68 percent of the time for Intermediate-term government bonds? What about 99.7 percent of the time?
Answer #1

68 percent of the time for Intermediate-term government bonds upper range
=5.4%+1*3.2%=8.60%

68 percent of the time for Intermediate-term government bonds lower range
=5.4%-1*3.2%=2.20%

so range of returns would you expect to see 68 percent is 2.20% to 8.60%

99.7 percent of the time for Intermediate-term government bonds upper range
=5.4%+3*3.2%=15.00%

99.7 percent of the time for Intermediate-term government bonds lower range
=5.4%-3*3.2%=-4.20%

so range of returns would you expect to see 99.7 percent is -4.20% to 15.00%